# Pair Production in time-dependent Electric field at Finite times

### v5

@@ IV.3.1 Momentum Spectra of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$
The longitudinal momentum significantly impacts the vacuum polarization function’s qualitative traits. To understand its dependence on $p_{\parallel}$, we comprehensively analyze the momentum spectra of both $u(p_{\parallel},t)$ and its associated counter term $v(p_{\parallel},t)$. Figure 14 shows the momentum spectra of $u(p_{\parallel},t)$ and $v(p_{\parallel},t).$ In the initial stages of particle formation, specifically at $t<0$, $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ display asymmetric Gaussian peaks with roughly similar profiles. During this time, $v(p_{\parallel},t)$ dominates over $u(p_{\parallel},t)$. Upon closer examination, it becomes evident that the peaks occur at different positions around $p_{\parallel}\approx 1$, as illustrated in Fig. 14(a). At $t=0$, the initial Gaussian-like structure of $u(p_{\parallel},t)$ is deformed, becoming bi-modal asymmetric structure. The spectrum $u(p_{\parallel},t)$ exhibits a peak at $p_{\parallel}\approx-0.35\ [m]$ and a dip at $p_{\parallel}\approx 0.35\ [m]$. This indicates a significant asymmetry in the function. On the other hand, the $v(p_{\parallel},t)$ spectrum maintains its Gaussian-like unimodal profile, with a peak at $p_{\parallel}=0$. During this stage, while the shape of $u(p_{\parallel},t)$ undergoes considerable changes and becomes asymmetric, the overall form of $v(p_{\parallel},t)$ remains unchanged. Due to the force factor ``$eE(t)"$ and corresponding longitudinal quasi-momentum $P(p_{\parallel},t)=(p_{\parallel}-eA(t))$, the spectrum moved to the left side of the origin, and its peak now located at $p_{\parallel}\approx-2[m]$ for $v(p_{\parallel},t)$ whereas small dip for $u(p_{\parallel},t)$ with some disruption in the tail ( $-1<p_{\parallel}<1$) as shown in the Figs.14(c-d) near to the transient stage.
@@
At $t=24\ [m^{-1}]$, we observe two distinct structures in the polarization and depolarization functions. The first structure appears at $p_{\parallel}=-2\ [m]$ due to earlier particle creation and annihilation events. The second structure, within $-1<p_{\parallel}<1$, exhibits varying amplitude oscillations, with the maximum occurring at zero longitudinal momentum. At this stage, $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ compete, as shown in Fig.14(d) and (e).
@@ IV.4 Dependence on Transverse momentum
In this subsection, we extensively investigate the effect of transverse momentum on the LMS. By plotting the time evolution of the LMS for different fixed values of $p_{\perp}$, the result is displayed in Figure 16. The spectra exhibit a smooth unimodal structure during the initial stages, as depicted in Fig. 16(a). Notably, the peak’s location shifts, as illustrated in Fig.16(b), accordingly with the $(p_{\parallel}-eA(t))=0$ as explained in section IV.2. Within a narrow spectral region for $-1<p_{\parallel}<1,$ some irregular oscillating structure is observed.
@@
One interesting fact observed was that the central peak magnitude diminished very rapidly compared to other peaks, and that oscillating structure does not exist for high transverse momentum at this time. The existence of oscillation in the central peak depends on the time taken for particles to reach the on-shell configuration from the initial off-shell state, a process governed by both transverse and longitudinal momenta (see Sec.IV.1).
@@
          $p_{\perp}[m]$           $t_{cp}[m^{-1}]$           $t_{sep}[m^{-1}]$           $t_{dis.}[m^{-1}]$
@@
Table 3: Effect of transverse mode on different time scale. The time scale $t_{cp}$ appearance of a central peak, $t_{sep}[m^{-1}]$ when two peaks become distinctly separated, and $t_{dis.}$ disappearance of the oscillation in LMS.
@@ IV.5 Transverse Momentum Spectrum
The effect of created pairs tends to have a peak in the momentum distribution function at non-zero transverse momentum, as discussed by Krajewska et al. Krajewska and Kamiński (2019). It is demonstrated that the transverse momentum distribution of real pairs exhibits an off-axis maximum above the one-photon threshold frequency (or high frequency); otherwise, the maximum for the transverse distribution function occurs at $p_{\perp}=0[m]$ in the presence of an oscillating time-dependent electric field. This study gives the clue that our case of off-axis maxima or multi-modal structure may also be observed due to multiphoton transition processes. To confirm this, we analyze the probability of pair production by calculating the amplitude in the two-level equivalent formalism Krajewska and Kamiński (2019). We decompose the time-dependent electric field $E(t)$ via the Fourier transformation into cosine and sine components. This decomposition isolates individual frequency components of the field, allowing us to study their effects on pair production at finite times. For a given frequency $\nu$, the corresponding electric field amplitude $\tilde{E}(\nu)$ facilitates the evaluation of the particle production spectrum at a specific finite time $t$ (see section A.2). The spectrum exhibits multiple peaks, forming a multimodal profile. Each peak corresponds to distinct energy or momentum exchanges between the field and the particle pairs, characteristic of multi-photon absorption. By considering different frequencies $\nu$ and their respective field amplitudes, we compute the pair production probability for each case. The total probability is obtained by summing the probabilities across different frequencies and then squaring the sum to derive the final result. The resulting spectrum features distinct peaks at various transverse momenta $p_{\perp}$. The peaks appear at non-zero transverse momenta, with their height and position dependent on the specific parameters of the electric field and the timing of the spectrum evaluation (see appendix A.2). The Fourier decomposition approach not only elucidates the origin of multiple peaks in the momentum spectrum but also establishes a connection between these features and the multi-photon process.
@@
Next, we investigate the impact of longitudinal momentum on TMS at finite time. We observe the interesting feature that different longitudinal values manifest distinct substructures in the spectrum at a given time, as explicitly verified from Figs. 18(a)-(b). The most common feature observed is that the magnitude of the distribution function decreases for higher longitudinal momentum. Before particles reach the on-shell condition, the spectra of created particles at times $t>3\tau$ exhibit a dramatic change in structure, contingent on the value of $p_{\parallel}.$ For $p_{\parallel}=0.1[m]$ and $p_{\parallel}=0.3[m]$, the maximum peak value occurs at vanishing transverse momentum. In contrast, for $p_{\parallel}=0.5[m]$, the TMS peak shifts to $p_{\perp}\approx 0.6[m]$ (Fig. 18(b)), indicating energy sharing between modes. At this finite time, we observe the sharing of longitudinal and transverse motion. The qualitative change in the shape of the spectrum, due to the influence of longitudinal momentum. Specifically, it requires more time for the quasi-particle to evolve into a real particle with the inclusion of longitudinal momentum when analyzing the TMS of quasi-particles at finite time. In this sense, we can say that the motion of the longitudinal and transverse directions is strongly coupled, and its impact is easily seen in Figs.18. These concerns can be understood by examining the highly oscillating phase kernel $\cos{(\Theta_{\bm{p}}(t_{0},t))}$ present in the quantum kinetic equation (QKE) Banerjee and Singh (2019). This kernel depends on the quasi-energy $\omega(\bm{p},t)$, which connects the longitudinal and transverse motions of the created particles. At the REPP stage, the substructure in the spectrum disappears, and the peak height for higher longitudinal momentum value is consistently lower, shown in Fig. 18(c).
@@ IV.5.1 Approximate expression for transverse momentum distribution function
$\mathrm{C}_{1}(p_{\perp},y)$ $=-\frac{4y|\Gamma_{1}\overline{\Gamma_{2}}|E_{0}\tau(1+p_{\perp}^{2})}{\sqrt{1+p_{\perp}^{2}+E_{0}^{2}\tau^{2}}}\cos{(\Upsilon)},$
@@
$-8|\Gamma_{1}\overline{\Gamma_{2}}|E_{0}\tau\cos{(\Upsilon)}\Biggl{(}\frac{(1+p_{\perp}^{2})(1+p_{\perp}^{2}+E_{0}\tau\sqrt{1+p_{\perp}^{2}+E_{0}^{2}\tau^{2}})}{(1+p_{\perp}^{2}+E_{0}^{2}\tau^{2})^{3/2}}$
@@
×$\times\frac{2yE_{0}\left(1+E_{0}\tau(E_{0}\tau+\sqrt{1+E_{0}^{2}\tau^{2}})\right)}{(1+E_{0}^{2}\tau^{2})^{3/2}}.$ (130)
@@
It shows that the behavior is symmetric about the origin, which can be understood as the presence of only the $p_{\perp}^{2}-$ term and the absence of the linear term in $p_{\perp}$, in comparison to the longitudinal momentum case as discussed in section IV.2.1. One more thing we notice is that at $p_{\perp}=0$, the behavior of the $\cos{(\Upsilon)}$ function changes drastically with time due to the presence of the time-dependent term $\ln{(1-y)}$. The oscillation frequency also depends on $p_{\perp}^{2}$, which means the frequency increases as the transverse momentum value increases. However, the coefficient of $\cos{(\Upsilon)}$ function plays a role here, giving a Gaussian envelope to the cosine function as discussed previously section IV.2.1. Within the width of the Gaussian profile, the oscillations survive and are seen in the spectrum.
@@
The above Eq. (135) clearly shows that in the REPP stage, the transverse distribution function has an invariable Gauss-like distribution whose half-width is defined by the field strength,$E_{0}$, and pulse duration $\tau.$
V Conclusion

### v6

@@ IV.3.1 Momentum Spectra of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$
The longitudinal momentum significantly impacts the vacuum polarization function’s qualitative traits. To understand its dependence on $p_{\parallel}$, we comprehensively analyze the momentum spectra of both $u(p_{\parallel},t)$ and its associated counter term $v(p_{\parallel},t)$. Figure 14 shows the momentum spectra of $u(p_{\parallel},t)$ and $v(p_{\parallel},t).$ In the initial stages of particle formation, specifically at $t<0$, $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ display asymmetric Gaussian peaks with roughly similar profiles. During this time, $v(p_{\parallel},t)$ dominates over $u(p_{\parallel},t)$. Upon closer examination, it becomes evident that the peaks occur at different positions around $p_{\parallel}\approx 1$, as illustrated in Fig. 14(a). At $t=0$, the initial Gaussian-like structure of $u(p_{\parallel},t)$ is deformed, becoming bi-modal asymmetric structure. The spectrum $u(p_{\parallel},t)$ exhibits a peak at $p_{\parallel}\approx-0.35\ [m]$ and a dip at $p_{\parallel}\approx 0.35\ [m]$. This indicates a significant asymmetry in the function. On the other hand, the $v(p_{\parallel},t)$ spectrum maintains its Gaussian-like unimodal profile, with a peak at $p_{\parallel}=0$. During this stage, while the shape of $u(p_{\parallel},t)$ undergoes considerable changes and becomes asymmetric, the overall shape of $v(p_{\parallel},t)$ remains unchanged. Due to the force factor eE(t)``eE(t)" and corresponding longitudinal quasi-momentum $P(p_{\parallel},t)=(p_{\parallel}-eA(t))$, the spectrum moved to the left side of the origin, and its peak now located at $p_{\parallel}\approx-2[m]$ for $v(p_{\parallel},t)$ whereas small dip for $u(p_{\parallel},t)$ with some disruption in the tail ( $1<p_{\parallel}<1$) as shown in the Figs.14(c-d) near to the transient stage.
@@
At $t=24\ [m^{-1}]$, we observe two distinct structures in the polarization and depolarization functions. The first structure appears at $p_{\parallel}=-2\ [m]$ due to earlier particle creation and annihilation events. The second structure, within $1<p_{\parallel}<1$, exhibits varying amplitude oscillations, with the maximum occurring at zero longitudinal momentum. At this stage, $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ compete, as shown in Fig.14(d) and (e).
@@ IV.4 Dependence on Transverse momentum
In this subsection, we extensively investigate the effect of transverse momentum on the LMS. By plotting the time evolution of the LMS for different fixed values of $p_{\perp}$, the result is displayed in Figure 16. The spectra exhibit a smooth unimodal structure during the initial stages, as depicted in Fig. 16(a). Notably, the peak’s location shifts, as illustrated in Fig.16(b), accordingly with the $(p_{\parallel}-eA(t))=0$ as explained in section IV.2. Within a narrow spectral region for $1<p_{\parallel}<1,$ some irregular oscillating structure is observed.
@@
One interesting fact observed was that the central peak magnitude diminished very rapidly compared to other peaks, and that oscillatory structure does not exist for high transverse momentum at this time. The existence of oscillation in the central peak depends on the time taken for particles to reach the on-shell configuration from the initial off-shell state, a process governed by both transverse and longitudinal momenta (see Sec.IV.1).
@@
   $p_{\perp}\penalty 10000\ [m]$    $t_{cp}\penalty 10000\ [m^{-1}]$    $t_{sep}\penalty 10000\ [m^{-1}]$    $t_{dis.}\penalty 10000\ [m^{-1}]$
@@
Table 3: Effect of transverse momentum $p_{\perp}$ on the time scales. $t_{cp}$ denotes the appearance of a central peak, $t_{sep}$ the time when two peaks become distinctly separated, and $t_{dis.}$ the disappearance of oscillations in LMS.
@@ IV.5 Transverse Momentum Spectrum
The effect of created pairs tends to have a peak in the momentum distribution function at non-zero transverse momentum, as discussed by Krajewska et al. Krajewska and Kamiński (2019). It is demonstrated that the transverse momentum distribution of real pairs exhibits an off-axis maximum above the one-photon threshold frequency (or high frequency); otherwise, the maximum for the transverse distribution function occurs at $p_{\perp}=0[m]$ in the presence of an oscillating time-dependent electric field. This study gives the clue that our case of off-axis maxima or multi-modal structure may also be observed due to multiphoton transition processes. To confirm this, we analyze the probability of pair production by calculating the amplitude in the two-level equivalent formalism Di Piazza et al. (2001); Krajewska and Kamiński (2019). We decompose the time-dependent electric field $E(t)$ via the Fourier transformation into cosine and sine components. This decomposition isolates individual frequency components of the field, allowing us to study their effects on pair production at finite times. For a given frequency $\nu$, the corresponding electric field amplitude $\tilde{E}(\nu)$ facilitates the evaluation of the particle production spectrum at a specific finite time $t$ (see appendix B). The spectrum exhibits multiple peaks, forming a multimodal profile. Each peak corresponds to distinct energy or momentum exchanges between the field and the particle pairs, characteristic of multi-photon absorption. By considering different frequencies $\nu$ and their respective field amplitudes, we compute the pair production probability for each case. The total probability is obtained by summing the probabilities across different frequencies and then squaring the sum to derive the final result. The resulting spectrum features distinct peaks at various transverse momenta $p_{\perp}$. The peaks appear at non-zero transverse momenta, with their height and position dependent on the specific parameters of the electric field and the timing of the spectrum evaluation (see appendix B). The Fourier decomposition approach not only elucidates the origin of multiple peaks in the momentum spectrum but also establishes a connection between these features and the multi-photon process.
@@
Next, we investigate the impact of longitudinal momentum on TMS at finite time. We observe the interesting feature that different longitudinal values manifest distinct substructures in the spectrum at a given time, as explicitly verified from Figs. 18(a)-(b). The most common feature observed is that the magnitude of the distribution function decreases for higher longitudinal momentum. Before particles reach the on-shell condition, the spectra of created particles at times $t>3\tau$ exhibit a dramatic change in structure, contingent on the value of $p_{\parallel}.$ For $p_{\parallel}=0.1[m]$ and $p_{\parallel}=0.3[m]$, the maximum peak value occurs at vanishing transverse momentum. In contrast, for $p_{\parallel}=0.5[m]$, the TMS peak shifts to $p_{\perp}\approx 0.6[m]$ (Fig. 18(b)), indicating energy sharing between modes. At this finite time, we observe a sharing of longitudinal and transverse motion, indicating that the two modes do not evolve independently. A similar interplay between longitudinal and transverse degrees of freedom has been reported in Ref. Majczak et al. (2024); Bechler et al. (2023).
The qualitative change in the shape of the spectrum, due to the influence of longitudinal momentum. Specifically, it requires more time for the quasi-particle to evolve into a real particle with the inclusion of longitudinal momentum when analyzing the TMS of quasi-particles at finite time. In this sense, we can say that the motion of the longitudinal and transverse directions is strongly coupled, and its impact is easily seen in Figs.18. These concerns can be understood by examining the highly oscillating phase kernel $\cos{(\Theta_{\bm{p}}(t_{0},t))}$ present in QKE Banerjee and Singh (2019). This kernel depends on the quasi-energy $\omega(\bm{p},t)$, which connects the longitudinal and transverse motions of the created particles. At the REPP stage, the substructure in the spectrum disappears, and the peak height for higher longitudinal momentum value is consistently lower, shown in Fig. 18(c).
@@ IV.5.1 Approximate expression for transverse momentum distribution function
$\mathrm{C}_{1}(p_{\perp},y)$ $=-\frac{4y|\Gamma_{1}\Gamma^{*}_{2}|E_{0}\tau(1+p_{\perp}^{2})}{\sqrt{1+p_{\perp}^{2}+E_{0}^{2}\tau^{2}}}\cos{(\Upsilon)},$
@@
$-8|\Gamma_{1}\Gamma^{*}_{2}|E_{0}\tau\cos{(\Upsilon)}\Biggl(\frac{(1+p_{\perp}^{2})(1+p_{\perp}^{2}+E_{0}\tau\sqrt{1+p_{\perp}^{2}+E_{0}^{2}\tau^{2}})}{(1+p_{\perp}^{2}+E_{0}^{2}\tau^{2})^{3/2}}$
@@
$\times\frac{2yE_{0}\left(1+E_{0}\tau(E_{0}\tau+\sqrt{1+E_{0}^{2}\tau^{2}})\right)}{(1+E_{0}^{2}\tau^{2})^{3/2}}.$ (130)
@@
It shows that the behavior is symmetric about the origin, which can be understood as the presence of only the $p_{\perp}^{2}$- term and the absence of the linear term in $p_{\perp}$, in comparison to the longitudinal momentum case as discussed in section IV.2.1. One more thing we notice is that at $p_{\perp}=0$, the behavior of the $\cos{(\Upsilon)}$ function changes drastically with time due to the presence of the time-dependent term $\ln{(1-y)}$. The oscillation frequency also depends on $p_{\perp}^{2}$, which means the frequency increases as the transverse momentum value increases. However, the coefficient of $\cos{(\Upsilon)}$ function plays a role here, giving a Gaussian envelope to the cosine function as discussed previously in section IV.2.1. Within the width of the Gaussian profile, the oscillations survive and are seen in the spectrum.
@@
[…]
V Possible physical interpretation
The quasiparticle distribution function $f(\bm{p},t)$ analyzed in previous sections provides a consistent dynamical description of pair production. However, its physical interpretation at finite times—when the external electric field is still active—requires careful consideration. Because the quasiparticle basis itself is time-dependent, $f(\bm{p},t)$ is basis-dependent and cannot be directly identified with the number of real, asymptotic particles Ilderton (2022); Dabrowski and Dunne (2016).
To establish a clear physical interpretation, we adopt a complementary Dirac sea picture (equivalent scattering) approach Avetissian et al. (2002). Following Ref. Ilderton (2022), we consider an electric field that acts only until a finite time $T$ and is then switched off abruptly. Within this framework, we solve the Dirac equation as a scattering problem: an initial negative-energy state evolves under the time-dependent field and, after switch-off, becomes a superposition of positive- and negative-energy free states Avetissian (2016). The transition amplitude to positive-energy states directly yields the probability $\mathcal{W}(\bm{p},T)$ for creating real electron–positron pairs.
Our goal is to prove the key identity:
$f(\bm{p},T)=\mathcal{W}(\bm{p},T),$ (136)
where $\mathcal{W}(\bm{p},T)$ is the probability of producing real electron–positron pairs if the field is switched off at time $T$. This equivalence bridges the formal quasiparticle description with an operationally measurable quantity.
V.1 Setup: Sauter pulse with finite duration
We modify the Sauter-pulse field so that it is nonzero only for $t<T$:
$E(t)=\begin{cases}E_{0}\operatorname{sech}^{2}\left(\frac{t}{\tau}\right),&-\infty<t<T,\\[4.0pt] 0,&t>T,\end{cases}$ (137)
with the corresponding vector potential
$A(t)=\begin{cases}-E_{0}\tau\tanh\left(\frac{t}{\tau}\right),&t<T,\\[4.0pt] -E_{0}\tau\tanh\left(\frac{T}{\tau}\right),&t>T.\end{cases}$ (138)
The problem naturally divides into three time intervals:
  • Region I ($t\to$-$\infty$): field off, initial negative-energy free state.
  • Region II ($\infty<t<T$): field on, exact Sauter-pulse solution applies.
  • Region III ($t>T$): field off, free-particle solutions.
V.1.1 Free Dirac spinors in Regions I and III
In the absence of an electric field ($E(t)=0$), the Dirac equation admits plane-wave solutions. For a particle with canonical momentum $\bm{p}=(p_{1},p_{2},p_{\parallel})$ and kinetic momentum $\Pi_{3}=p_{\parallel}-eA$ (where $A$ is constant), the energy is $\mathcal{E}=\sqrt{m^{2}+p_{\perp}^{2}+\Pi_{3}^{2}}$. The positive- and negative-energy spinors for spin $r=1$ are (see Appendix C for explicit forms and conventions):
$\xi_{\bm{p},1}^{(+)}(t)=\frac{1}{\sqrt{2\mathcal{E}(\mathcal{E}-\Pi_{3})}}\begin{pmatrix}m\\ 0\\ -\mathcal{E}+\Pi_{3}\\ p_{1}+ip_{2}\end{pmatrix}e^{-i\mathcal{E}t},\qquad\xi_{\bm{p},1}^{(-)}(t)=\frac{1}{\sqrt{2\mathcal{E}(\mathcal{E}+\Pi_{3})}}\begin{pmatrix}m\\ 0\\ \mathcal{E}+\Pi_{3}\\ p_{1}+ip_{2}\end{pmatrix}e^{i\mathcal{E}t}.$ (139)
These satisfy the orthonormality condition $\xi_{\bm{p},1}^{(\lambda)}(t)]^{\dagger}\xi_{\bm{p},1}^{(\lambda^{\prime})}(t)=\delta_{\lambda\lambda^{\prime}}$ with respect to the Dirac inner product Schweber (1961); Avetissian (2016).
In Region I ($t\to$-$\infty$):
$\Pi_{3}=P_{0}\equiv p_{\parallel}-eA(-\infty)=p_{\parallel}-E_{0}\tau,\quad\mathcal{E}=\omega_{0}\equiv\sqrt{m^{2}+p_{\perp}^{2}+P_{0}^{2}}.$
In Region III ($t>T$):
$\Pi_{3}=p_{\parallel}-eA(T),\quad\mathcal{E}=\sqrt{m^{2}+p_{\perp}^{2}+\Pi_{3}^{2}}.$
V.1.2 Exact solution in Region II
In Region II, the field is the full Sauter pulse. The mode function satisfies
$\Big(\partial_{t}^{2}+iE(t)+\omega^{2}(\bm{p},t)\Big)\psi_{\bm{p}}(t)=0.$ (140)
Introducing the variable
$y=\frac{1}{2}\left[1+\tanh\left(\frac{t}{\tau}\right)\right],\quad y\in[0,Y],\quad Y=y(T),$ (141)
this equation transforms into a hypergeometric differential equation. The general solution regular at $y=0$ is
$\psi_{\bm{p}}^{(+)}(y)=\mathcal{A}\;y^{-i\tau\omega_{0}/2}(1-y)^{i\tau\omega_{1}/2}\,{}_{2}F_{1}(a,b;c;y),$ (142)
where
$\omega_{0}$ $=\sqrt{\epsilon_{\perp}^{2}(p_{\perp})+P_{0}^{2}},\quad\omega_{1}=\sqrt{\epsilon_{\perp}^{2}(p_{\perp})+P_{1}^{2}},$
and $\mathcal{A}$ is a normalization constant.
Matching to Region I as $t\to$-$\infty$ fixes
$\mathcal{A}=\frac{1}{\sqrt{2\omega_{0}(\omega_{0}-P_{0})}}.$
Thus, the normalized positive-energy mode function in Region II is
$\psi_{\bm{p}}^{(+)}(y)=\frac{y^{-i\tau\omega_{0}/2}(1-y)^{i\tau\omega_{1}/2}}{\sqrt{2\omega_{0}(\omega_{0}-P_{0})}}\;{}_{2}F_{1}(a,b;c;y).$ (143)
The corresponding Dirac spinor for spin $r=1$ in Region II is (see Appendix C for derivation):
$\Psi_{\bm{p},1}^{(+)}(t)=\begin{pmatrix}m\\ 0\\ i\partial_{t}+P(p_{\parallel},t)\\ p_{1}+ip_{2}\end{pmatrix}\psi_{\bm{p}}^{(+)}(t),\qquad P(p_{\parallel},t)=p_{\parallel}-eA(t).$ (144)
At $t=T$, the solution from Region II must match continuously onto a superposition of free states in Region III:
$\Psi_{\bm{p},1}^{(+)}(T)=\mathcal{R}(\bm{p})\,\xi_{\bm{p},1}^{(-)}(T)+\mathcal{I}(\bm{p})\,\xi_{\bm{p},1}^{(+)}(T),$ (145)
where $\mathcal{R}(\bm{p})$ is the pair production amplitude (reflection from the Dirac sea) and $\mathcal{I}(\bm{p})$ is the transmission amplitude Avetissian (2016).
Projecting both sides onto $\xi_{\bm{p},1}^{(-)}(T)$ using orthonormality yields:
$\mathcal{R}(\bm{p})=\left[\xi_{\bm{p},1}^{(-)}(T)\right]^{\dagger}\Psi_{\bm{p},1}^{(+)}(T).$ (146)
Substituting the explicit forms (139) and (144) and simplifying (see Appendix C for details) gives:
$\mathcal{R}(\bm{p})=-i\sqrt{\frac{\mathcal{E}+\Pi_{3}}{2\mathcal{E}}}\Big[\partial_{t}\psi_{\bm{p}}^{(+)}(T)+i\mathcal{E}\,\psi_{\bm{p}}^{(+)}(T)\Big].$ (147)
The probability to create an electron–positron pair with momentum $\bm{p}$ and a given spin is $|\mathcal{R}(\bm{p})|^{2}$. Summing over the two independent spin states gives a factor of 2, so the total pair production probability is:
$\mathcal{W}(\bm{p},T)=2|\mathcal{R}(\bm{p})|^{2}.$ (148)
Inserting the expression for $\psi_{\bm{p}}^{(+)}(t)$ from Eq. (143) and writing derivatives in terms of the $y$-variable leads to:
$\mathcal{W}(\bm{p},T)=\frac{\mathcal{E}+\Pi_{3}}{2\mathcal{E}\,\omega_{0}(\omega_{0}-P_{0})}\left|\frac{2}{\tau}Y(1-Y)\frac{ab}{c}f_{1}+i\big(\mathcal{E}-(1-Y)\omega_{0}-Y\omega_{1}\big)f_{2}\right|^{2},$ (149)
where $Y=y(T)$, $f_{1}={}_{2}F_{1}(a+1,b+1;c+1;Y)$, and $f_{2}={}_{2}F_{1}(a,b;c;Y)$.
A direct term-by-term comparison reveals that Eq. (149) is identical to the quasiparticle distribution function $f(\bm{p},y)$ given in Eq. (72), evaluated at $y=Y$. Specifically, the prefactor matches $|N^{(+)}(\bm{p})|^{2}\left(1+\frac{P}{\omega}\right)$ at $y=Y$, and the expression inside the absolute value is exactly the same as in Eq. (72).
Therefore,
$\boxed{\,f(\bm{p},T)=\mathcal{W}(\bm{p},T)\,}$ (150)
where $f(\bm{p},T)\equiv f(\bm{p},y(T))$.
The physical significance of the quasiparticle distribution function within the adiabatic basis—arising from the Bogoliubov transformation—is that it acquires the interpretation of a real particle number if the external electric field is switched off at a finite time $T$. This occurs because the Bogoliubov transformation captures the essential mixing between negative- and positive-energy states induced by the time-dependent electric field. Once the field is terminated, this mixing freezes, and the quasiparticle excitations evolve into free asymptotic particles.
While an abrupt switch-off represents an idealization, previous studies have shown that the same physical conclusions are obtained when the external field decays rapidly but smoothly (Ref. Álvarez-Domínguez et al. (2023)). This ensures that the identification $f(\bm{p},T)=\mathcal{W}(\bm{p},T)$ provides a physically meaningful and robust interpretation of the finite-time quasiparticle distribution.
Equation (150) establishes that the quasiparticle distribution function at a finite time $T$ equals the probability to produce real, observable electron–positron pairs if the electric field is switched off at that instant. This result justifies the detailed analysis of $f(\bm{p},t)$ presented in Section IV: the oscillatory structures in the momentum spectra and their dependence on longitudinal and transverse momentum reflect the dynamical build-up of real electron–positron pairs, not merely transient quasiparticle features.
Moreover, by linking $f(\bm{p},t)$ to the probability of observable pair production in a scenario where the electric field is switched off at a finite time, the formalism becomes directly relevant to pulsed-field experiments, such as those anticipated at next-generation high-intensity laser facilities.
VI Conclusion